Nabil is considering buying a house while he is at university. The house costs 110,000 dollars today. Renting out part of the house and living in the rest over his five years at school will net, after expenses, 1,250 dollars per month. He estimates that he will sell the house after five years for 119,000 dollars. If Nabil's MARR is 18%, compounded monthly, should he buy the house? Use present worth analysis {Perform all calculations using 5 significant figures and round any monetary answers to the nearest cent}.

a.) What is the present worth of this project?

b.) Should Nabil buy the house?
Answer #1

I = 18% = 18% / 12 = 1.5% per month

t = 5 yrs = 5*12 = 60 months

NPW = -110000 + 1250 * (P/A, 1.5%,60) + 119000 * (P/F, 1.5%,60)

= -110000 + 1250 * 39.380269 + 119000 * 0.409296

= -12068.44

As NPW is negative, he should not buy the house